OpenStudy (anonymous):

hey can u guys help me out with this one 100ml of soft drink A contains 75g sugar and weighs 110g . Calculate the density of the soft drink

OpenStudy (anonymous):

d = m/V => d = 110/75 = 1,467g/ml

OpenStudy (anonymous):

if density is mass over volume , the volume cannot divide the mass , but the opposite ... my confusion is more about the masses 75 g and 110 g ....

OpenStudy (anonymous):

the question is density of the soft drink not the sugar.. so density of softdrink = mass of soft drink / volume of soft drink.. you dont need to use mass of sugar..